# Approximation in Lattice Field Theories

### v1

@@ 5 Conclusion
$10^{2}$ $10^{3}$ $10^{6}$ $10^{2}$ $10^{3}$ $10^{6}$
1 $7.83\times 10^{-6}$ $8.62\times 10^{-3}$ $1.47\times 10^{-3}$ $5.90\times 10^{-3}$ $9.08\times 10^{-3}$ $1.44\times 10^{-3}$
2 $1.43\times 10^{-2}$ $2.87\times 10^{-3}$ $2.34\times 10^{-5}$ $1.86\times 10^{-2}$ $3.97\times 10^{-3}$ $4.79\times 10^{-4}$
3 $8.74\times 10^{-4}$ $1.07\times 10^{-4}$ $7.86\times 10^{-3}$ $6.35\times 10^{-3}$ $1.32\times 10^{-3}$ $7.86\times 10^{-3}$
4 $1.54\times 10^{-2}$ $5.71\times 10^{-3}$ $2.70\times 10^{-4}$ $2.03\times 10^{-2}$ $6.84\times 10^{-3}$ $2.10\times 10^{-4}$
5 $6.86\times 10^{-4}$ $1.33\times 10^{-2}$ $1.76\times 10^{-3}$ $6.79\times 10^{-3}$ $1.38\times 10^{-2}$ $1.96\times 10^{-3}$
6 $1.24\times 10^{-2}$ $5.26\times 10^{-3}$ $5.29\times 10^{-3}$ $1.72\times 10^{-2}$ $6.38\times 10^{-3}$ $5.45\times 10^{-3}$
7 $8.90\times 10^{-3}$ $2.17\times 10^{-5}$ $1.23\times 10^{-2}$ $38\times 10^{-2} 1.1$$7\times 10^{-3} 1.1$$19\times 10^{-2}$
8 $1.39\times 10^{-3}$ $1.31\times 10^{-3}$ $2.17\times 10^{-3}$ $6.73\times 10^{-3}$ $2.48\times 10^{-3}$ $2.41\times 10^{-3}$
9 $6.08\times 10^{-4}$ $7.77\times 10^{-3}$ $4.52\times 10^{-3}$ $4.79\times 10^{-3}$ $8.18\times 10^{-3}$ $4.25\times 10^{-3}$
11 $1.095$ $1.079$ $1.102$ $1.32\times 10^{-2}$ $1.57\times 10^{-3}$ $1.91\times 10^{-2}$
21 $1.40\times 10^{-2}$ $1.21\times 10^{-2}$ $9.76\times 10^{-3}$ $3.03\times 10^{-3}$ $07\times 10^{-3} 1.1$$14\times 10^{-3}$
22 $1.081$ $1.079$ $1.078$ $5.19\times 10^{-4}$ $2.40\times 10^{-3}$ $3.23\times 10^{-3}$
31 $1.18\times 10^{-2}$ $23\times 10^{-2} 1.1$$14\times 10^{-2}$ $5.09\times 10^{-4}$ $1.40\times 10^{-3}$ $1.37\times 10^{-4}$
32 $1.18\times 10^{-2}$ $1.23\times 10^{-2}$ $1.40\times 10^{-2}$ $5.78\times 10^{-4}$ $7.66\times 10^{-4}$ $2.80\times 10^{-3}$
33 $1.080$ $1.078$ $1.074$ $1.58\times 10^{-3}$ $2.77\times 10^{-3}$ $6.32\times 10^{-3}$
41 $1.05\times 10^{-2}$ $5.76\times 10^{-3}$ $7.34\times 10^{-3}$ $4.09\times 10^{-4}$ $5.09\times 10^{-3}$ $3.55\times 10^{-3}$
42 $4.16\times 10^{-2}$ $4.56\times 10^{-2}$ $4.66\times 10^{-2}$ $4.68\times 10^{-3}$ $9.96\times 10^{-4}$ $2.71\times 10^{-4}$
43 $4.47\times 10^{-2}$ $4.60\times 10^{-2}$ $5.24\times 10^{-2}$ $1.60\times 10^{-3}$ $7.90\times 10^{-5}$ $5.80\times 10^{-3}$
44 $1.072$ $1.076$ $1.079$ $8.87\times 10^{-3}$ $4.16\times 10^{-3}$ $2.55\times 10^{-3}$
51 $4.65\times 10^{-2}$ $4.34\times 10^{-2}$ $5.25\times 10^{-2}$ $2.44\times 10^{-5}$ $2.85\times 10^{-3}$ $5.19\times 10^{-3}$
52 $1.51\times 10^{-2}$ $1.20\times 10^{-2}$ $1.03\times 10^{-2}$ $3.63\times 10^{-3}$ $6.91\times 10^{-4}$ $6.88\times 10^{-4}$
53 $5.91\times 10^{-2}$ $3.87\times 10^{-2}$ $3.82\times 10^{-2}$ $1.20\times 10^{-2}$ $6.75\times 10^{-3}$ $7.28\times 10^{-3}$
54 $1.57\times 10^{-2}$ $1.09\times 10^{-2}$ $3.10\times 10^{-3}$ $4.66\times 10^{-3}$ $3.18\times 10^{-4}$ $7.35\times 10^{-3}$
55 $1.084$ $1.083$ $1.095$ $1.63\times 10^{-3}$ $2.24\times 10^{-3}$ $1.30\times 10^{-2}$
61 $4.76\times 10^{-2}$ $4.61\times 10^{-2}$ $5.14\times 10^{-2}$ $3.95\times 10^{-4}$ $3.23\times 10^{-4}$ $3.98\times 10^{-3}$
62 $5.18\times 10^{-2}$ $4.58\times 10^{-2}$ $4.77\times 10^{-2}$ $4.72\times 10^{-3}$ $9.18\times 10^{-4}$ $1.37\times 10^{-3}$
63 $3.67\times 10^{-3}$ $8.59\times 10^{-3}$ $1.42\times 10^{-2}$ $6.83\times 10^{-3}$ $2.16\times 10^{-3}$ $2.84\times 10^{-3}$
64 $1.06\times 10^{-2}$ $8.54\times 10^{-3}$ $1.46\times 10^{-2}$ $1.98\times 10^{-4}$ $2.10\times 10^{-3}$ $3.18\times 10^{-3}$
65 $1.69\times 10^{-2}$ $7.12\times 10^{-3}$ $1.94\times 10^{-2}$ $5.13\times 10^{-3}$ $3.80\times 10^{-3}$ $7.76\times 10^{-3}$
66 $1.062$ $1.082$ $1.087$ $84\times 10^{-2} 1.1$$17\times 10^{-3}$ $5.76\times 10^{-3}$
71 $1.09\times 10^{-2}$ $1.21\times 10^{-2}$ $7.89\times 10^{-3}$ $4.45\times 10^{-4}$ $1.07\times 10^{-3}$ $3.36\times 10^{-3}$
72 $4.52\times 10^{-2}$ $5.08\times 10^{-2}$ $4.61\times 10^{-2}$ $1.59\times 10^{-3}$ $4.14\times 10^{-3}$ $1.85\times 10^{-4}$
73 $4.23\times 10^{-2}$ $15\times 10^{-2} 4.4$$40\times 10^{-2}$ $4.04\times 10^{-3}$ $4.04\times 10^{-3}$ $1.91\times 10^{-3}$
74 $6.53\times 10^{-3}$ $1.22\times 10^{-2}$ $1.32\times 10^{-2}$ $4.17\times 10^{-3}$ $1.29\times 10^{-3}$ $1.96\times 10^{-3}$
75 $4.54\times 10^{-2}$ $5.11\times 10^{-2}$ $4.45\times 10^{-2}$ $1.95\times 10^{-3}$ $4.70\times 10^{-3}$ $2.18\times 10^{-3}$
76 $5.28\times 10^{-2}$ $4.22\times 10^{-2}$ $4.64\times 10^{-2}$ $6.41\times 10^{-3}$ $3.56\times 10^{-3}$ $7.19\times 10^{-4}$
77 $1.091$ $1.078$ $1.074$ $8.93\times 10^{-3}$ $3.25\times 10^{-3}$ $6.60\times 10^{-3}$
81 $4.85\times 10^{-2}$ $4.72\times 10^{-2}$ $4.98\times 10^{-2}$ $1.36\times 10^{-3}$ $1.02\times 10^{-3}$ $2.42\times 10^{-3}$
82 $1.43\times 10^{-2}$ $1.93\times 10^{-2}$ $1.21\times 10^{-2}$ $2.75\times 10^{-3}$ $7.63\times 10^{-3}$ $9.33\times 10^{-4}$
83 $4.42\times 10^{-2}$ $4.82\times 10^{-2}$ $5.09\times 10^{-2}$ $1.76\times 10^{-3}$ $1.96\times 10^{-3}$ $4.19\times 10^{-3}$
84 $5.00\times 10^{-2}$ $4.73\times 10^{-2}$ $5.50\times 10^{-2}$ $3.73\times 10^{-3}$ $1.37\times 10^{-3}$ $7.97\times 10^{-3}$
85 $1.17\times 10^{-2}$ $1.34\times 10^{-2}$ $8.52\times 10^{-3}$ $3.50\times 10^{-4}$ $2.72\times 10^{-3}$ $2.31\times 10^{-3}$
86 $4.26\times 10^{-2}$ $3.79\times 10^{-2}$ $4.65\times 10^{-2}$ $2.75\times 10^{-3}$ $7.62\times 10^{-3}$ $5.71\times 10^{-4}$
87 $7.06\times 10^{-3}$ $1.58\times 10^{-2}$ $8.42\times 10^{-3}$ $3.73\times 10^{-3}$ $4.70\times 10^{-3}$ $2.23\times 10^{-3}$
88 $1.071$ $1.103$ $1.086$ $9.34\times 10^{-3}$ $2.08\times 10^{-2}$ $3.86\times 10^{-3}$
91 $3.38\times 10^{-2}$ $4.90\times 10^{-2}$ $4.39\times 10^{-2}$ $1.25\times 10^{-2}$ $2.75\times 10^{-3}$ $3.09\times 10^{-3}$
92 $4.37\times 10^{-2}$ $4.54\times 10^{-2}$ $4.68\times 10^{-2}$ $2.71\times 10^{-3}$ $9.32\times 10^{-4}$ $6.27\times 10^{-4}$
93 $1.19\times 10^{-2}$ $7.39\times 10^{-3}$ $8.69\times 10^{-3}$ $2.99\times 10^{-4}$ $3.13\times 10^{-3}$ $2.18\times 10^{-3}$
94 $4.37\times 10^{-2}$ $3.96\times 10^{-2}$ $4.07\times 10^{-2}$ $2.24\times 10^{-3}$ $5.89\times 10^{-3}$ $4.87\times 10^{-3}$
95 $4.45\times 10^{-2}$ $4.67\times 10^{-2}$ $4.85\times 10^{-2}$ $2.35\times 10^{-3}$ $3.26\times 10^{-4}$ $1.67\times 10^{-3}$
96 $1.38\times 10^{-2}$ $6.85\times 10^{-3}$ $1.03\times 10^{-2}$ $2.16\times 10^{-3}$ $3.54\times 10^{-3}$ $1.38\times 10^{-3}$
97 $8.79\times 10^{-3}$ $1.04\times 10^{-2}$ $1.33\times 10^{-2}$ $1.97\times 10^{-3}$ $9.22\times 10^{-4}$ $2.37\times 10^{-3}$
98 $1.76\times 10^{-2}$ $1.54\times 10^{-2}$ $7.38\times 10^{-3}$ $5.76\times 10^{-3}$ $3.55\times 10^{-3}$ $3.79\times 10^{-3}$
99 $1.083$ $1.094$ $1.077$ $2.21\times 10^{-3}$ $1.20\times 10^{-2}$ $3.44\times 10^{-3}$
Table A1: Table of numerically evaluated 1 and 2-pt correlation functions for $M=9$ and $N=10^{2},10^{3},$ and $10^{6}$, along with the right hand side expressions of eqns. (4) and (5) with these values plugged in. Index $k$ corresponds to $\langle x(k)\rangle$ in the first three columns and eqn. (4) with $i=k$. Similarly, index klkl corresponds to $\langle x(k)x(l)\rangle$ in the first three columns and eqn. (4) with $i=k$ and $t=l$. It is easy to that the 1-pt functions and the equation results, on average, grow closer to 0 for larger $N$. The behavior of the 2-pt functions also gets closer to values from eqn. (17), but the diagonal terms have more error due to larger spacing in the Markov Chain construction.

### v2

@@ 5 Conclusion
$40$ $100$ $1000$ $40$ $100$ $1000$
11 $2.14\times 10^{-2}$ $7.57\times 10^{-4}$ $3.97\times 10^{-3}$ $7.38\times 10^{-3}$ $7.86\times 10^{-3}$ $1.15\times 10^{-3}$
22 $9.51\times 10^{-3}$ $1.87\times 10^{-2}$ $2.06\times 10^{-3}$ $4.36\times 10^{-3}$ $9.84\times 10^{-3}$ $4.70\times 10^{-3}$
33 $1.38\times 10^{-2}$ $2.79\times 10^{-3}$ $5.82\times 10^{-3}$ $2.78\times 10^{-4}$ $1.04\times 10^{-2}$ $7.75\times 10^{-3}$
44 $1.58\times 10^{-2}$ $6.43\times 10^{-3}$ $1.35\times 10^{-5}$ $2.29\times 10^{-3}$ $1.36\times 10^{-3}$ $2.00\times 10^{-3}$
55 $1.29\times 10^{-2}$ $1.05\times 10^{-2}$ $1.48\times 10^{-5}$ $1.30\times 10^{-3}$ $2.47\times 10^{-3}$ $2.41\times 10^{-3}$
66 $1.56\times 10^{-2}$ $2.52\times 10^{-3}$ $9.86\times 10^{-3}$ $1.21\times 10^{-3}$ $1.02\times 10^{-2}$ $6.86\times 10^{-3}$
77 $2.40\times 10^{-2}$ $1.53\times 10^{-2}$ $5.34\times 10^{-3}$ $9.55\times 10^{-3}$ $6.67\times 10^{-3}$ $2.77\times 10^{-3}$
88 $8.87\times 10^{-3}$ $2.90\times 10^{-3}$ $1.45\times 10^{-3}$ $4.94\times 10^{-3}$ $9.66\times 10^{-3}$ $3.80\times 10^{-3}$
99 $7.09\times 10^{-3}$ $3.51\times 10^{-3}$ $3.69\times 10^{-3}$ $6.53\times 10^{-3}$ $4.73\times 10^{-3}$ $1.23\times 10^{-3}$
$11$ $1.085$ $1.082$ $1.092$ $1.18\times 10^{-3}$ $1.81\times 10^{-5}$ $1.04\times 10^{-2}$
2121 $1.34\times 10^{-2}$ $1.39\times 10^{-2}$ $4.74\times 10^{-3}$ $1.94\times 10^{-3}$ $3.28\times 10^{-3}$ $5.39\times 10^{-3}$
$22$ $1.091$ $1.076$ $1.073$ $6.70\times 10^{-3}$ $5.43\times 10^{-3}$ $7.43\times 10^{-3}$
$31$ $2.02\times 10^{-2}$ $1.31\times 10^{-2}$ $7.97\times 10^{-3}$ $8.60\times 10^{-3}$ $2.34\times 10^{-3}$ $2.54\times 10^{-3}$
$32$ $1.19\times 10^{-2}$ $5.79\times 10^{-3}$ $1.30\times 10^{-2}$ $6.45\times 10^{-4}$ $4.58\times 10^{-3}$ $2.19\times 10^{-3}$
$33$ $1.075$ $1.075$ $1.077$ $7.67\times 10^{-3}$ $6.29\times 10^{-3}$ $3.92\times 10^{-3}$
$41$ $1.34\times 10^{-2}$ $4.85\times 10^{-3}$ $2.94\times 10^{-3}$ $1.81\times 10^{-3}$ $5.59\times 10^{-3}$ $7.06\times 10^{-3}$
$42$ $5.13\times 10^{-2}$ $4.10\times 10^{-2}$ $4.78\times 10^{-2}$ $3.48\times 10^{-3}$ $5.13\times 10^{-3}$ $1.62\times 10^{-3}$
$43$ $4.61\times 10^{-2}$ $3.99\times 10^{-2}$ $4.67\times 10^{-2}$ $1.04\times 10^{-3}$ $5.47\times 10^{-3}$ $3.37\times 10^{-4}$
$44$ $1.068$ $1.076$ $1.072$ $1.47\times 10^{-2}$ $4.72\times 10^{-3}$ $9.08\times 10^{-3}$
$51$ $46\times 10^{-2} 4.4$$44\times 10^{-2}$ $5.19\times 10^{-2}$ $2.55\times 10^{-3}$ $1.68\times 10^{-3}$ $5.12\times 10^{-3}$
$52$ $1.54\times 10^{-3}$ $1.56\times 10^{-2}$ $4.68\times 10^{-3}$ $9.15\times 10^{-3}$ $4.19\times 10^{-3}$ $5.79\times 10^{-3}$
$53$ $4.31\times 10^{-2}$ $3.51\times 10^{-2}$ $4.00\times 10^{-2}$ $3.83\times 10^{-3}$ $1.05\times 10^{-2}$ $5.77\times 10^{-3}$
$54$ $1.39\times 10^{-2}$ $3.03\times 10^{-3}$ $2.08\times 10^{-2}$ $2.61\times 10^{-3}$ $6.32\times 10^{-3}$ $9.41\times 10^{-3}$
$55$ $1.088$ $1.071$ $1.076$ $4.21\times 10^{-3}$ $1.05\times 10^{-2}$ $4.82\times 10^{-3}$
$61$ $5.11\times 10^{-2}$ $4.75\times 10^{-2}$ $4.57\times 10^{-2}$ $3.78\times 10^{-3}$ $1.03\times 10^{-3}$ $6.64\times 10^{-4}$
$62$ $4.62\times 10^{-2}$ $4.21\times 10^{-2}$ $4.65\times 10^{-2}$ $1.24\times 10^{-3}$ $3.81\times 10^{-3}$ $7.25\times 10^{-4}$
$63$ $1.52\times 10^{-2}$ $5.54\times 10^{-3}$ $4.19\times 10^{-3}$ $3.28\times 10^{-3}$ $5.12\times 10^{-3}$ $6.46\times 10^{-3}$
$64$ $82\times 10^{-2} 1.1$$16\times 10^{-2}$ $1.58\times 10^{-2}$ $6.54\times 10^{-3}$ $1.13\times 10^{-3}$ $4.72\times 10^{-3}$
$65$ $1.56\times 10^{-2}$ $1.28\times 10^{-2}$ $9.25\times 10^{-3}$ $4.49\times 10^{-3}$ $1.70\times 10^{-3}$ $1.28\times 10^{-3}$
$66$ $1.079$ $1.081$ $1.086$ $5.39\times 10^{-3}$ $4.93\times 10^{-4}$ $4.14\times 10^{-3}$
$71$ $7.82\times 10^{-3}$ $7.67\times 10^{-3}$ $8.72\times 10^{-3}$ $4.05\times 10^{-3}$ $3.24\times 10^{-3}$ $1.98\times 10^{-3}$
$72$ $4.40\times 10^{-2}$ $4.03\times 10^{-2}$ $4.59\times 10^{-2}$ $2.97\times 10^{-3}$ $5.77\times 10^{-3}$ $1.11\times 10^{-5}$
$73$ $4.24\times 10^{-2}$ $5.33\times 10^{-2}$ $4.92\times 10^{-2}$ $4.39\times 10^{-3}$ $7.40\times 10^{-3}$ $3.23\times 10^{-3}$
$74$ $6.30\times 10^{-3}$ $1.81\times 10^{-3}$ $1.08\times 10^{-2}$ $5.31\times 10^{-3}$ $7.93\times 10^{-3}$ $7.41\times 10^{-4}$
$75$ $4.67\times 10^{-2}$ $4.23\times 10^{-2}$ $4.09\times 10^{-2}$ $3.93\times 10^{-4}$ $3.61\times 10^{-3}$ $4.46\times 10^{-3}$
$76$ $5.37\times 10^{-2}$ $3.69\times 10^{-2}$ $4.99\times 10^{-2}$ $5.69\times 10^{-3}$ $8.72\times 10^{-3}$ $3.33\times 10^{-3}$
$77$ $1.094$ $1.080$ $1.080$ $9.50\times 10^{-3}$ $1.48\times 10^{-3}$ $1.48\times 10^{-3}$
$81$ $3.75\times 10^{-2}$ $4.49\times 10^{-2}$ $4.38\times 10^{-2}$ $9.50\times 10^{-3}$ $1.42\times 10^{-3}$ $2.37\times 10^{-3}$
$82$ $1.53\times 10^{-2}$ $1.07\times 10^{-2}$ $8.62\times 10^{-3}$ $3.55\times 10^{-3}$ $1.65\times 10^{-4}$ $2.07\times 10^{-3}$
$83$ $4.91\times 10^{-2}$ $4.57\times 10^{-2}$ $5.03\times 10^{-2}$ $1.75\times 10^{-3}$ $1.21\times 10^{-4}$ $4.14\times 10^{-3}$
$84$ $4.14\times 10^{-2}$ $5.39\times 10^{-2}$ $4.63\times 10^{-2}$ $5.05\times 10^{-3}$ $7.65\times 10^{-3}$ $3.55\times 10^{-4}$
$85$ $8.15\times 10^{-3}$ $1.22\times 10^{-2}$ $7.98\times 10^{-3}$ $2.91\times 10^{-3}$ $1.77\times 10^{-3}$ $2.99\times 10^{-3}$
$86$ $58\times 10^{-2} 4.4$$45\times 10^{-2}$ $4.33\times 10^{-2}$ $1.61\times 10^{-3}$ $1.69\times 10^{-3}$ $2.94\times 10^{-3}$
$87$ $9.96\times 10^{-3}$ $8.50\times 10^{-3}$ $53\times 10^{-2} 1.1$$14\times 10^{-3}$ $1.72\times 10^{-3}$ $3.94\times 10^{-3}$
$88$ $1.070$ $1.082$ $1.110$ $1.19\times 10^{-2}$ $1.65\times 10^{-4}$ $2.65\times 10^{-2}$
$91$ $4.21\times 10^{-2}$ $3.77\times 10^{-2}$ $4.39\times 10^{-2}$ $4.82\times 10^{-3}$ $8.08\times 10^{-3}$ $2.39\times 10^{-3}$
$92$ $4.66\times 10^{-2}$ $5.14\times 10^{-2}$ $4.96\times 10^{-2}$ $5.34\times 10^{-4}$ $4.55\times 10^{-3}$ $3.70\times 10^{-3}$
$93$ $1.57\times 10^{-2}$ $5.96\times 10^{-3}$ $1.14\times 10^{-2}$ $3.96\times 10^{-3}$ $3.95\times 10^{-3}$ $7.09\times 10^{-4}$
$94$ $4.52\times 10^{-2}$ $3.51\times 10^{-2}$ $4.68\times 10^{-2}$ $1.47\times 10^{-3}$ $9.74\times 10^{-3}$ $6.39\times 10^{-4}$
$95$ $4.38\times 10^{-2}$ $5.76\times 10^{-2}$ $4.54\times 10^{-2}$ $2.84\times 10^{-3}$ $1.08\times 10^{-2}$ $9.10\times 10^{-4}$
$96$ $1.22\times 10^{-2}$ $07\times 10^{-2} 1.1$$13\times 10^{-2}$ $2.96\times 10^{-4}$ $2.33\times 10^{-4}$ $1.43\times 10^{-4}$
$97$ $1.35\times 10^{-2}$ $6.35\times 10^{-3}$ $5.09\times 10^{-3}$ $2.34\times 10^{-3}$ $3.56\times 10^{-3}$ $5.26\times 10^{-3}$
$98$ $8.70\times 10^{-3}$ $24\times 10^{-2} 1.1$$16\times 10^{-2}$ $1.78\times 10^{-3}$ $9.93\times 10^{-4}$ $7.06\times 10^{-4}$
$99$ $1.097$ $1.081$ $1.072$ $1.28\times 10^{-2}$ $6.97\times 10^{-4}$ $8.57\times 10^{-3}$